# Rolled Gaussian process models for curves on manifolds

### v1

@@ 10.6 Proofs
For every $t$, suppose $F(y):=E\{\rho^{2}(y,x(t))\}<\infty$ at a point $y$, and hence on all $M$. For part $(i)$, when every point $p\in M$ has empty cut locus the function $y\mapsto F(y)$ is geodesically convex on $M$ with a global minimizer that is the unique stationary point, and the Fréchet mean of $x(t)$ is hence unique (e.g. Afsari, 2011). On such $M$, $\exp_{\gamma(t)}:T_{\gamma(t)}M\to M$ is bijective, and it is thus sufficient to show that $E\{\exp^{-1}_{\gamma(t)}(x(t))\}=0$ (e.g. Le, 2001). Since $\exp^{-1}_{\gamma(t)}(x(t))=P^{\gamma}_{t\leftarrow 0}P^{c}_{1\leftarrow 0}\{y(t)-\tilde{m}(t)\},$ and the parallel transports are linear isometries, the result follows since $E\{y(t)-\tilde{m}(t)\}=0$.
@@
since $M$ is symmetric and $\sigma_{\gamma(t)}$ is an isometry on $M$. Let $v\in T_{\gamma(t)}M$ be such that $\exp_{\gamma(t)}(v)=x(t)$. Since the distribution of $x(t)$ is assumed to possess even symmetry, it is of the form $(\exp_{\gamma(t)})_{\#}\lambda$ of a distribution $\lambda$ with Lebesgue-density $f$ that is symmetric around the origin in $T_{\gamma(t)}M$ such that $f(v)=f(-v)$; this holds regardless of whether $v$ is unique when $x(t)\notin\mathcal{C}(\gamma(t))$ or a measurable selection is made from the set of all $v$ such that $\exp_{\gamma(t)}(v)=x(t)$. Since the isometry $\sigma_{\gamma(t)}$ reverses geodesics through ${\gamma(t)}$, we have the relation $\exp_{\gamma(t)}(-v)=\sigma_{\gamma(t)}(\exp_{\gamma(t)}(v))$. Therefore,
@@
10.6.2 Proof of Proposition 3.1
For every $t$, note that $y_{b}(t)\sim N_{d}(\tilde{m}(t),\tilde{k}(t,t))$. For every $t$, the orthonormal frame $E^{\gamma}(t):T_{\gamma(0)}M\to T_{\gamma(t)}M$ is a linear operator. Consequently,
$V^{\gamma}(t)\sim N_{d}\left(E^{\gamma}(t)\tilde{m}(t),E^{\gamma}(t)\tilde{k}(t,t)E^{\gamma}(t)^{\top}\right)$
with values in the vector space $(T_{\gamma(t)}M,\langle\cdot,\cdot\rangle_{\gamma(t)})$, where $E^{\gamma}(t)^{\top}$ is the corresponding adjoint operator. Consider for distinct times $t,t^{\prime}$ the random vector $(V^{\gamma}(t),V^{\gamma}(t^{\prime}))$ with values in the direct sum inner product space $\left(T_{\gamma(t)}M\oplus T_{\gamma(t^{\prime})}M{,}\langle\cdot,\cdot\rangle_{\gamma(t)}+\langle\cdot,\cdot\rangle_{\gamma(t^{\prime})}\right)$. From the argument above it is clear that $(V^{\gamma}(t),V^{\gamma}(t^{\prime}))$ is Gaussian, and that its mean is $(E^{\gamma}(t)\tilde{m}(t),E^{\gamma}(t^{\prime})\tilde{m}(t^{\prime}))$. From Appendix 10.1 we observe that the covariance between $V^{\gamma}(t)$ and $V^{\gamma}(t^{\prime})$ is given by linear operator $\Omega_{tt^{\prime}}:T_{\gamma(t)}M\oplus T_{\gamma(t^{\prime})}M\to T_{\gamma(t)}M\oplus T_{\gamma(t^{\prime})}M$ defined by
@@
where $\Omega_{ab}:=E^{\gamma}(a)\tilde{k}(a,b)E^{\gamma}(b)^{\top}$. The discussion above holds for any curve $\gamma$, and thus for $\gamma=\tilde{m}^{\uparrow}_{b}$.
@@
It is now straightforward to extend the above to the case of $r$ times $t_{1},\ldots,t_{r}$ involving a direct sum of $r$ tangent spaces and corresponding $r\times r$ block covariance operator comprising ($r\choose 2$ linear operators $\Omega_{t_{i}t_{j}}$ for $i,j=1,\ldots,r$. This establishes the existence of finite-dimensional marginal Gaussian distributions corresponding to the choices $m$ and $k$. If the family of finite-dimensional marginal distributions constitutes a projective family, existence and uniqueness of a Gaussian vector field $t\mapsto V^{\gamma}(t)$ along $\gamma$ follows from Kolmogorov’s extension theorem (e.g., Bhattacharya and Waymire, 2007). Proof of projectivity follows straightforwardly from Proposition 16 in Hutchinson et al. (2021), who consider general Gaussian vector fields not necessarily along curves $\gamma$.
10.6.3 Proof of Proposition 3.2
@@
Given a mean function $m$ and the kernel $k$ for the $\mathbb{R}^{d}$-valued Gaussian process $z$, the pair $(b,U)$ determines $\tilde{m}=Um$ and $\tilde{k}=UkU^{\top}$ in $T_{b}M$, the mean and kernel, respectively, of the Gaussian process $y_{b}(t)$ in $T_{b}M$; the unique rolling of $\tilde{m}$ onto $M$ then results in $\tilde{m}^{\uparrow}_{b}$. Upon setting $\gamma=\tilde{m}^{\uparrow}_{b}$, the Gaussian vector field $V^{\gamma}(t)$ is obtained by action of the frame $E^{\gamma}(t)$, determined by $E^{\gamma}(0)=(P^{c}_{1\leftarrow 0}u_{1},\ldots,P^{c}_{1\leftarrow 0}u_{d})$, on the Gaussian process $y_{b}(t)$; in other words, on $T_{b}M$, the map $(U,m,k)\to V^{\gamma}$ is injective. The curve $x$ on $M$ is then defined as $x(t)=\exp_{\gamma(t)}(V^{\gamma}(t))$.

### v2

@@ Appendix D Proofs of main results
For every $t$, suppose $F(y):=E\{\rho^{2}(y,x(t))\}<\infty$ at a point $y$, and hence on all $M$. For part $(i)$, when every point $p\in M$ has empty cut locus the function $y\mapsto F(y)$ is geodesically convex on $M$ with a global minimizer that is the unique stationary point, and the Fréchet mean of $x(t)$ is hence unique (e.g. Afsari, 2011). On such $M$, $\exp_{\gamma(t)}:T_{\gamma(t)}M\to M$ is bijective, and it is thus sufficient to show that $E\{\exp^{-1}_{\gamma(t)}(x(t))\}=0$ (e.g. Le, 2001). Since $\exp^{-1}_{\gamma(t)}(x(t))=P^{\gamma}_{t\leftarrow 0}P^{c}_{1\leftarrow 0}\{y(t)-m_{b}(t)\},$ and the parallel transports are linear isometries, the result follows since $E\{y(t)-m_{b}(t)\}=0$.
@@
since $M$ is symmetric and $\sigma_{\gamma(t)}$ is an isometry on $M$. Let $v\in T_{\gamma(t)}M$ be such that $\exp_{\gamma(t)}(v)=x(t)$. The distribution of $x(t)$ is of the form $(\exp_{\gamma(t)})_{\#}\lambda$ of a distribution $\lambda$ with Lebesgue-density $f$ that is symmetric around the origin in $T_{\gamma(t)}M$ such that f(v)=f(v)f(v)=f(-v); this holds regardless of whether $v$ is unique when $x(t)\notin\mathcal{C}(\gamma(t))$ or a measurable selection is made from the set of all $v$ such that $\exp_{\gamma(t)}(v)=x(t)$. As a consequence, the distribution of $x(t)$ possesses geodesic symmetry. Since the isometry $\sigma_{\gamma(t)}$ reverses geodesics through $\gamma(t)$}, we have the relation $\exp_{\gamma(t)}(-v)=\sigma_{\gamma(t)}(\exp_{\gamma(t)}(v))$. Therefore,
@@
D.2 Proof of Proposition 4.1
$E^{\gamma}=\{e_{1},\ldots,e_{d}\}$ be a frame along $\gamma$, defined in Appendix A.2. For every $t$, note that $y_{b}(t)-m_{b}(t)\sim N_{d}(0,K_{b}(t,t))$. For every $t$, the orthonormal frame $E^{\gamma}(t):T_{\gamma(0)}M\to T_{\gamma(t)}M$ is a linear operator. Consequently,
$V^{\gamma}(t)\sim N_{d}\left(0,E^{\gamma}(t)K_{b}(t,t)E^{\gamma}(t)^{\top}\right)$
with values in the vector space $(T_{\gamma(t)}M,\langle\cdot,\cdot\rangle_{\gamma(t)})$, where $E^{\gamma}(t)^{\top}$ is the corresponding adjoint operator. Consider for distinct times $t,t^{\prime}$ the random vector $(V^{\gamma}(t),V^{\gamma}(t^{\prime}))$ with values in the direct sum inner product space $\left(T_{\gamma(t)}M\oplus T_{\gamma(t^{\prime})}M{,}\langle\cdot,\cdot\rangle_{\gamma(t)}+\langle\cdot,\cdot\rangle_{\gamma(t^{\prime})}\right)$. Then, $(V^{\gamma}(t),V^{\gamma}(t^{\prime}))$ is a zero-mean Gaussian with covariance
@@
where $\Omega_{ab}:=E^{\gamma}(a)K_{b}(a,b)E^{\gamma}(b)^{\top}$ (see Appendix A.1) The discussion above holds for any curve $\gamma$, and thus for $\gamma=m_{b}^{\uparrow}$.
@@
It is now straightforward to extend the above to the case of $r$ times $t_{1},\ldots,t_{r}$ involving a direct sum of $r$ tangent spaces and corresponding $r\times r$ block covariance operator comprising $r\choose 2$ linear operators $\Omega_{t_{i}t_{j}}$ for $i,j=1,\ldots,r$. This establishes the existence of finite-dimensional marginal Gaussian distributions corresponding to the choices $m$ and $k$. If the family of finite-dimensional marginal distributions constitutes a projective family, existence and uniqueness of a Gaussian vector field $t\mapsto V^{\gamma}(t)$ along $\gamma$ follows from Kolmogorov’s extension theorem (e.g., Bhattacharya and Waymire, 2007). Proof of projectivity follows straightforwardly from Proposition 16 in Hutchinson et al. (2021), who consider general Gaussian vector fields not necessarily along curves $\gamma$.
D.3 Proof of Proposition 4.2
@@
Given a mean $m$ and the covariance $K$ for the $\mathbb{R}^{d}$-valued Gaussian process $z$, the pair $(b,U)$ determines $m_{b}=Um$ and $K_{b}=UKU^{\top}$ in $T_{b}M$, the mean and kernel, respectively, of the Gaussian process $y_{b}(t)$ in $T_{b}M$; the unique rolling of $m_{b}$ onto $M$ then results in $m_{b}^{\uparrow}$. Upon setting $\gamma=m_{b}^{\uparrow}$, the Gaussian vector field $V^{\gamma}(t)$ is obtained by action of the frame $E^{\gamma}(t)$, determined by $E^{\gamma}(0)=(P^{c}_{1\leftarrow 0}u_{1},\ldots,P^{c}_{1\leftarrow 0}u_{d})$, on the Gaussian process $y_{b}(t)$; in other words, the map $(U,m,K)\to V^{\gamma}$ is injective. The curve $x$ on $M$ is then defined as $x(t)=\exp_{\gamma(t)}(V^{\gamma}(t))$.